# Using Diffusion Models to do Data Assimilation

### v1

@@ 2.3 Diffusion-based data assimilation
This type of conditional image generation using diffusion modeling is precisely what is required for DA. The typical approach is to reuse the tried-and-true recipe from generating images, i.e., we train a diffusion model to take in noise and then generate (atmospheric) system states. The only difference to the diffusion model outlined in Section 2.2.2 is that we condition on the current set of observations, $\mathbf{y}_{k}$, i.e., the denoiser is the minimizer of the loss function
$\ell=E_{\mathbf{x}_{k},\mathbf{y}_{k}\sim p(\mathbf{x}_{k},\mathbf{y}_{k})}E_{n\sim\mathcal{N}(0,t^{2})}||\mathbf{x}_{k}-\mathbf{D}(\mathbf{v}=\mathbf{x}_{k}+n,\mathbf{y}_{k},t)||_{2}^{2},$ (13)
@@
where $p(\mathbf{x}_{k},\mathbf{y}_{k})$ is the joint posterior and this conditioning is well-understood in the diffusion modeling literature (see, e.g., Batzolis et al., 2021; Chung et al., 2023; Qu et al., 2024).
@@
Note that the procedure goes like this: we collect a large set of training data in the form of a time-series of system states $\mathbf{x}_{k}$ and observations $\mathbf{y}_{k}$. We subsequently set up a denoiser in the form of a neural network to minimize the loss function in (13). This training is once and for all and (usually) never repeated. The result is a diffusion model that takes in the latest observations and then generates an atmospheric system state. Examples of the use of diffusion models that work in this way include Rozet and Louppe (2023); Chung et al. (2023); Manshausen et al. (2024); Pathak et al. (2024); Qu et al. (2024); Li et al. (2025), but the list of papers is rapidly expanding.
We will show below that we can interpret the output of such a diffusion model as samples from the following pdf:
$p(\mathbf{x}_{k}|\mathbf{y}_{k})\propto p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k}),$ (14)
where the training set corresponds to samples from the prior, $p(\mathbf{x}_{k})$, i.e., the training set is a long time-series of past weather. Note that the above posterior distribution is different from the posterior distribution in (1) used in conventional DA. The key difference lies in the prior. For diffusion DA, the prior is implicitly defined by the training dataset, which is not updated from cycle-to-cycle and, thus, will be referred to here as a climatological prior. This implies that this climatological prior is independent of observations from the recent past. This form of DA with a climatological prior is unusual from the perspective of conventional DA because conventional DA has always valued the cycling nature of the DA problem and has therefore traditionally focused on the posterior distribution (1) using a cycling prior.
Other variants of diffusion DA supplement the observations $\mathbf{y}_{k}$ with an additional predictor in the form of a forecast, $\mathbf{f}_{k}$, either produced by the diffusion DA or by other means (see, e.g., Huang et al., 2024). We will show below that since the prior corresponds to the training set (which is again not cycled in these works), the forecast is implicitly treated like an additional observation. Hence, the targeted posterior distribution in these works is
$p(\mathbf{x}_{k}|\mathbf{y}_{k},\mathbf{f}_{k})\propto p(\mathbf{y}_{k},\mathbf{f}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k}).$ (15)
The forecast is a function of all past observations, but the prior is not updated and remains climatological (assuming no re-training at each DA cycle). Instead, the likelihood is modified to incorporate the forecast as a kind of additional observation. Such DA systems are thus somewhat in between a cycling, conventional DA system and the diffusion DA described just above with a climatological prior. We will label DA systems that incorporate a forecast as an extended likelihood DA system, because the observations are extended to include the forecast and this extended set of observations are then used within a likelihood as in (15).
@@ 2.4 Other forms of ML assisted data assimilation
Our focus here is on diffusion DA and on answering questions related to how generative AI may be able to replace an entire ensemble DA system and what it might mean if it does. But ML methods have other uses in DA which we want to briefly acknowledge. Perhaps most importantly, there is a flurry of recent activity on using, for example, the ERA5 reanalysis for training ML-based forecast models (see, e.g., Price et al., 2025; Lam et al., 2023; Kochkov et al., 2024; Li et al., 2024; Mardani et al., 2025). In terms of probability distributions, these ML methods sample conditional distributions of the type
$p(\mathbf{x}_{k+T}|\mathbf{x}_{k},\mathbf{x}_{k-1}),$ (16)
where $T$ is the desired forecast lead time and where we (arbitrarily) stopped the conditioning two time steps backwards in time (as, e.g., GenCast does, Price et al. (2025)). In terms of DA, these ML-based forecast models can be used as the model in these systems, but then the conventional DA system is still required. In this sense, ML-forecasting as part of a cycling DA system is very different from the diffusion DA systems we study here, because we have chosen to only consider diffusion models that aim at replacing the DA system, i.e., retain the forecast model unchanged. Experiments with replacing the physics-based forecast model with ML type models within an ensemble DA system are currently ongoing (see, e.g., Adrian et al., 2025).
3 Diffusion modeling in a linear, stochastic dynamical system
We illustrate here how a diffusion DA system emulates various forms of Bayesian posterior distributions by considering a linear, stochastic dynamical system for which we can compute the various posterior distributions without approximation using the Kalman filter formalism. The model describes the time evolution of an $n_{x}$-dimensional state, $\mathbf{x}$, governed by the stochastic differential equation (SDE)
$\text{d}\mathbf{x}=-\frac{1}{2}\mathbf{x}\text{d}s+\text{d}\beta_{s}$ (17)
where $\beta_{s}$ is a standard Wiener process and $s$ denotes physical time. Observations $\mathbf{y}_{k}$ are collected $\Delta s$ time units apart via a linear observation operator, $\mathbf{H}$, generating $n_{y}$ observations at each observation time
$\mathbf{y}_{k}=\mathbf{H}\mathbf{x}_{k}+\boldsymbol{\varepsilon}_{k}.$ (18)
@@ 3 Diffusion modeling in a linear, stochastic dynamical system
The climatological prior for this problem is obtained from a long simulation of the dynamics,  (17). For a diffusion DA system, this long model run is used as the training set. One reason we chose this simple problem setup is that we know the climatological prior analytically. For the linear dynamics (17), we can compute the climatological prior by solving the corresponding steady-state Fokker-Planck equation
@@
Note that similar, uncorrelated Gaussians have been used extensively to study conventional DA systems (see, e.g., Snyder et al., 2008; Bickel et al., 2008; Bengtsson et al., 2008; Snyder et al., 2015; Morzfeld and Hodyss, 2023; Hodyss and Morzfeld, 2023).
3.1 DA with a climatological prior
We now sample the Bayesian posterior distribution with a climatological prior. We first consider the Kalman approach in order to compute the moments of the posterior distribution analytically. We then show that the typical training paradigm used in diffusion modeling leads to this same result.
@@ 3.1.1 Bayesian posterior
We can use the Kalman filter to compute the exact mean and covariance of the posterior distribution with a climatological prior (20). Since this posterior distribution is Gaussian, we only need to compute the posterior mean and posterior covariance. Using the fact that the climatological prior has mean zero and the identity covariance matrix, we find
$\overline{\mathbf{x}}^{a}_{k}$ $=\mathbf{H}^{T}[\mathbf{H}\mathbf{H}^{T}+r\mathbf{I}]^{-1}\mathbf{y}_{k},$ (21)
$\mathbf{P}^{a}$ $=\mathbf{I}_{N}-\mathbf{H}^{T}[\mathbf{H}\mathbf{H}^{T}+r\mathbf{I}]^{-1}\mathbf{H},$ (22)
for the posterior mean and covariance. Since we assume that $\mathbf{H}$ only contains a subset of the rows of the identity matrix, there is no correlation between the state variables and we can examine the posterior mean element-wise and consider only the diagonal elements of the posterior covariance matrix. For an observed grid-point, we have
$[\mathbf{x}^{a}_{k}]^{j}$ $=\frac{1}{1+r}y,$ (23)
@@ 3.1.2 Diffusion DA
backward in virtual time (from $t=T$ to $t=0$).
We use Tweedie’s formula
@@
to compute the score, but avoid neural networks. Rather, we use the fact that the conditional expectation in (26) is the minimum of
$\ell=E_{\mathbf{x}_{k},\mathbf{y}_{k}\sim p(\mathbf{x}_{k},\mathbf{y}_{k})}E_{n\sim\mathcal{N}(0,t^{2})}||\mathbf{x}_{k}-\mathbf{D}(\mathbf{v}=\mathbf{x}_{k}+n,t,\mathbf{y}_{k})||_{2}^{2},$ (27)
Note that we have
$p(\mathbf{x}_{k},\mathbf{y}_{k})=p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k})$ (28)
which represents a joint posterior distribution using a climatological pdf. This implies that the training set consists of samples from a long simulation of the dynamical system under consideration paired up with observations at each time in the training set.
The minimum of (27) for a linear, Gaussian problem is a kind of Kalman filter of the form
$\mathbf{D}(\mathbf{v},t,\mathbf{y}_{k})=E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k})}[\mathbf{x}_{k}]=\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}$ (29)
@@
$\mathbf{R}=\begin{bmatrix}t^{2}\mathbf{I}_{n_{x}}&\mathbf{0}\\ \mathbf{0}&r\mathbf{I}_{n_{o}}\end{bmatrix},\quad\hat{\mathbf{H}}=\begin{bmatrix}\mathbf{I}_{n_{x}}\\ \mathbf{H}\end{bmatrix},\quad\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\end{bmatrix}.$ (30)
Using the Kalman formalism here allows us to obtain analytical expressions for both the denoiser and the score. On the other hand, in practice one would obtain very similar results with a well-trained denoiser. In this sense, replacing the denoiser with equation (29) is equivalent to assuming that the diffusion model is trained on an essentially infinite training dataset.
The diffusion model thus becomes
$\text{d}\mathbf{v}=-\frac{2}{t}\left[\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-\mathbf{v}\right]\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}$ (31)
Note that the “effective” observation error covariance in (30) implies that, early in the reverse process ($t^{2}\gg r$), the state is drawn towards the observations, $\mathbf{y}_{k}$, but when $t^{2}\ll r$, the state is drawn towards $\mathbf{v}$. Similarly, note that the noise term vanishes as $t\rightarrow 0$ but the drift term takes on a greater significance. These two effects ensure that the ensemble obtained from the diffusion model has the correct mean and variance.
The simple observation operator we consider here implies that there is no covariance between state variables, and thus we can consider a single element of the vector $\mathbf{v}$, which we call $v$ for simplicity. For an observed element of $\mathbf{v}$ at the $i^{th}$ gridpoint we have
$\text{d}v=-\frac{2}{t}\Big{[}\frac{\frac{r}{1+r}}{\frac{r}{1+r}+t^{2}}v+\frac{\frac{t^{2}}{1+r}}{\frac{r}{1+r}+t^{2}}y-v\Big{]}\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}$ (32)
@@
$\begin{split}v(0)=\frac{\frac{r}{1+r}}{\frac{r}{1+r}+T^{2}}v(T)+\frac{r}{(1+r)^{2}}\frac{T^{2}}{\frac{r}{1+r}(\frac{r}{1+r}+T^{2})}y+\frac{r}{1+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\end{split}$ (33)
Note that $v(0)$ is the random variable of interest and we can compute its mean as
@@
$\langle v(0)\rangle_{T\rightarrow\infty}=\frac{1}{1+r}y$ (35)
which is the posterior mean we obtained via the Kalman filter in (23), i.e., without diffusion.
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{r}{1+r}}{\frac{r}{1+r}+T^{2}}(v(T)-\langle v(T)\rangle)^{2}+\Bigg{\langle}\Bigg{(}\frac{r}{1+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\end{split}$ (36)
where we have used the fact that the cross-term vanishes. The stochastic integral on the right-hand side requires the use $\hat{\textrm{o}}$ isometry, i.e.,
$\begin{split}\Bigg{\langle}\Bigg{(}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}=\Bigg{\langle}\Bigg{(}-\int_{0}^{T}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\\ =2\int_{0}^{T}\frac{t}{(\frac{r}{1+r}+t^{2})^{2}}\text{d}t=-2\int_{T}^{0}\frac{t}{(\frac{r}{1+r}+t^{2})^{2}}\text{d}t\end{split}$ (37)
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{r^{2}}{(1+r)^{2}}T^{2}}{(\frac{r}{1+r}+T^{2})^{2}}+\frac{r}{1+r}\frac{T^{2}}{\frac{r}{1+r}+T^{2}}\end{split}$ (38)
@@
which is the same variance we obtained via the Kalman filter in (24).
Hence, we have now shown that the use of a long time-series of samples from a dynamical system as a training set for a diffusion model results in a diffusion DA system that samples a Bayesian posterior distribution constructed with a climatological prior.
3.2 DA with a cycling prior
We now consider the more conventional DA approach of sampling the Bayesian posterior distribution (1), with a cycling prior that propagates information from one cycle to the next.

### v2

@@ 3.2 Diffusion data assimilation
A common extension of the diffusion modeling described just above is conditional image generation in which one generates images of a requested scene given text prompts (see, e.g., Zhan et al., 2024; Ding et al., 2025). Conditional image generation via diffusion modeling is applicable to DA: We train a diffusion model to take in noise and then generate (atmospheric) system states. The only difference to the diffusion model outlined above is that we condition on the observations, $\mathbf{y}_{k}$, i.e., the denoiser is the minimizer of the loss function
[…]
@@
where $p(\mathbf{x}_{k},\mathbf{y}_{k})$ is the joint posterior. Conditioning is well-understood in the diffusion modeling literature (see, e.g., Batzolis et al., 2021; Chung et al., 2023; Qu et al., 2024).
@@
Diffusion DA thus follows this recipe: (i) we collect a large training dataset in the form of a time-series of system states $\mathbf{x}_{k}$ and observations $\mathbf{y}_{k}$; (ii) we set up a denoiser in the form of a neural network to minimize the loss function in (18). This training is once and for all and (usually) never repeated. The result is a diffusion model that takes in the latest observations and then generates an atmospheric system state. Examples of the use of diffusion models that work in this way include Qu et al. (2024); Pathak et al. (2024); Li et al. (2025). Other variants of diffusion DA supplement the observations $\mathbf{y}_{k}$ with an additional predictor in the form of a forecast, either produced by the diffusion DA or by other means (see, e.g., Huang et al., 2024).
One can also use approximations to the denoiser that do not need to be explicitly trained on a particular observation network (Chung et al., 2023; Rozet and Louppe, 2023; Manshausen et al., 2024). The idea here is to think of the posterior PDF to be sampled from as
$p_{t}(\mathbf{v}|\mathbf{y})\propto p_{t}(\mathbf{y}|\mathbf{v})p_{t}(\mathbf{v})$ (19)
where we emphasize that $\mathbf{v}$ is a function of the time, $t$, of the diffusion process and we incorporate a subscript $t$ on the PDF here to emphasize that the PDF changes with the time of the diffusion process. Thus, as time in the diffusion process approaches zero, i.e., $t\rightarrow 0$, then $p_{t}(\mathbf{v})\rightarrow p_{0}(\mathbf{v})$, which is the PDF describing the training set. When using a long time series of past data for training, $p_{0}(\mathbf{v})$ is the “climatological” PDF. Perhaps more importantly, note further that $p_{t}(\mathbf{y}|\mathbf{v})$ approaches the known Gaussian observation likelihood also as $t\rightarrow 0$. We use this fact next.
The desired score function is
$\nabla\log(p_{t}(\mathbf{v}|\mathbf{y}))=\nabla\log(p_{t}(\mathbf{y}|\mathbf{v}))+\nabla\log(p_{t}(\mathbf{v}))$ (20)
where the gradient is over the state, $\mathbf{v}$. The second term on the right is the unconditional score found in (14) and determined from the climatological prior. The first term on the right is approximated as
$\nabla\log(p_{t}(\mathbf{y}|\mathbf{v}))\approx\nabla\log(p_{0}(\mathbf{y}|D(\mathbf{v},t)))$ (21)
where $p_{0}(\mathbf{y}|D)=p_{t\rightarrow 0}(\mathbf{y}|D)$, which as stated above is equal to the known Gaussian observation likelihood but conditioned on the result from the the unconditional denoiser, $D(\mathbf{v},\tau)$, determined from the climatological prior. Assuming the observation likelihood has a covariance matrix, $\mathbf{R}=r\mathbf{I}$, equation (20) can be written as
[…]
This approximation offers enormous computational savings as a new observation type can be assimilated without any re-training. On the other hand, the approximation in (21) causes errors which we will explore further below.
@@ 3.3 Other forms of ML assisted data assimilation
There has been a flurry of recent activity using, for example, the ERA5 reanalysis for training ML-based forecast models (see, e.g., Price et al., 2025; Lam et al., 2023; Kochkov et al., 2024; Li et al., 2024; Mardani et al., 2025). In terms of probability distributions, these ML methods sample conditional distributions of the type $p(\mathbf{x}_{k+T}|\mathbf{x}_{k},\mathbf{x}_{k-1})$, where $T$ is the desired forecast lead time and where we (arbitrarily) stopped the conditioning two time steps backwards in time (as, e.g., GenCast does, Price et al. (2025)). ML-based forecast models can replace or augment the physics-based forecast model in current DA systems, with some results reported by Adrian et al. (2025). Our focus, however, is find out if and how generative AI may replace an entire DA system, which is quite different from integrating data-driven forecast models into existing DA systems.
4 Analysis of diffusion DA in a linear Gaussian system
We show how different training datasets used in diffusion DA lead to the sampling of different Bayesian posterior distributions. Our analysis uses a linear, stochastic dynamical system, because it allows us to perform calculations without approximations. For example, we can directly compute the various posterior distributions. Moreover, a very similar problem setup has been used extensively to study ensemble DA systems and in particular the collapse of particle filters (see, e.g., Snyder et al., 2008; Bickel et al., 2008; Bengtsson et al., 2008; Snyder et al., 2015; Morzfeld and Hodyss, 2023; Hodyss and Morzfeld, 2023).
4.1 Problem setup
We consider the linear SDE
$\text{d}\mathbf{x}=-\frac{1}{2}\mathbf{x}\text{d}s+\text{d}\beta_{s},$ (23)
where $\mathbf{x}$ is the $n_{x}$-dimensional system state, $\beta_{s}$ is a Wiener process, and $s$ is physical time. Observations $\mathbf{y}_{k}$ are collected $\Delta s$ time units apart via a linear observation operator, $\mathbf{H}$, generating $n_{y}$ observations
$\mathbf{y}_{k}=\mathbf{H}\mathbf{x}_{k}+\boldsymbol{\varepsilon}_{k},$ (24)
@@ 4.1 Problem setup
The climatological prior for this problem is obtained from a long simulation of the dynamics (23). For a diffusion DA system, this long model run is used as the training set. The climatological prior, however, can also be calculated analytically by solving the steady-state Fokker-Planck equation
@@
Thus, the training dataset are snapshots of a long time series of the solution of (23) or, equivalently, draws from the standard Gaussian, i.e. a Gaussian with mean zero and covariance as the identity.
4.2 Diffusion DA and the climatological prior
We consider the Bayesian posterior distribution with a climatological prior. This distribution is Gaussian and, hence, determined by the mean and (co-)variance, which can be computed analytically. We show that a diffusion DA that is trained with a long simulation, generates samples from this Bayesian posterior distribution.
@@ 4.2.1 Bayesian posterior
The Bayesian posterior in this case is
$p(\mathbf{x}_{k}|\mathbf{y}_{k})\propto p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k})$ (27)
where the prior is the climatological one in equation (26) and the observation likelihood is
$p(\mathbf{y}_{k}|\mathbf{x}_{k})\propto\exp\left(-\frac{1}{2}(\mathbf{y}_{k}-\mathbf{H}\mathbf{x})^{T}\mathbf{R}^{-1}(\mathbf{y}_{k}-\mathbf{H}\mathbf{x})\right)$ (28)
where $\mathbf{R}=r\mathbf{I}_{n_{y}}$. Thus, the posterior mean and covariance for this posterior distribution can readily be computed by, for example, multiplying (26) with (28) and subsequently completing the square within the argument of the exponential, to obtain
$\overline{\mathbf{x}}^{a}_{k}$ $=\mathbf{H}^{T}[\mathbf{H}\mathbf{H}^{T}+r\mathbf{I}_{n_{y}}]^{-1}\mathbf{y}_{k},$ (29)
$\mathbf{P}^{a}$ $=\mathbf{I}_{n_{x}}-\mathbf{H}^{T}[\mathbf{H}\mathbf{H}^{T}+r\mathbf{I}_{n_{y}}]^{-1}\mathbf{H}.$ (30)
Since we assume that $\mathbf{H}$ only contains a subset of the rows of the identity matrix, all matrices that appear above are diagonal (no correlation between state variables). Thus, we can examine the posterior mean element-wise and consider only the diagonal elements of the posterior covariance matrix. For an observed grid-point, we have
$[\overline{\mathbf{x}}^{a}_{k}]^{j}$ $=\frac{1}{1+r}y,$ (31)
@@ 4.2.2 Diffusion DA (exact denoiser)
backward in virtual time (from $t=T$ to $t=0$). We use Tweedie’s formula
@@
to compute the score, but we avoid neural networks, and compute the conditional expectation (or denoiser) analytically, which is equivalent to assuming a well-trained neural network. Note that
$p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k})\propto p(\mathbf{v},\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k})\propto p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{v}|\mathbf{x}_{k})p(\mathbf{x}_{k}),$ (35)
because $\mathbf{v}$ and $\mathbf{y}_{k}$ are independent, $p(\mathbf{y}_{k}|\mathbf{x}_{k})$ is the familiar likelihood in equation (28), and $p(\mathbf{x}_{k})$ is the climatological prior (the long training dataset). Due to our simple definition of the forward process, the reverse process is such that
$\mathbf{v}=\mathbf{x}_{k}+\mathbf{n},\quad\mathbf{n}\sim\mathcal{N}(0,t^{2}\mathbf{I}),$ (36)
so that
$p(\mathbf{v}|\mathbf{x}_{k})\propto\exp\left(-\frac{1}{2t^{2}}(\mathbf{v}-\mathbf{x}_{k})^{T}(\mathbf{v}-\mathbf{x}_{k})\right).$ (37)
The likelihood is defined by (28) and, therefore, putting everything together gives
$p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k})\propto\exp\left(-\frac{1}{2}(\hat{\mathbf{y}}-\hat{\mathbf{H}}\mathbf{x}_{k})^{T}\hat{\mathbf{R}}^{-1}(\hat{\mathbf{y}}-\hat{\mathbf{H}}\mathbf{x}_{k})\right)\exp\left(-\frac{1}{2}\mathbf{x}_{k}^{T}\mathbf{x}_{k}\right),$ (38)
@@
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\end{bmatrix},\quad\hat{\mathbf{H}}=\begin{bmatrix}\mathbf{I}_{n_{x}}\\ \mathbf{H}\end{bmatrix},\quad\hat{\mathbf{R}}=\begin{bmatrix}t^{2}\mathbf{I}_{n_{x}}&\mathbf{0}\\ \mathbf{0}&r\mathbf{I}_{n_{o}}\end{bmatrix}$. (39)
Again, completing the square within the exponential finds the formula for the conditional mean
$E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k})}[\mathbf{x}_{k}]=\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\hat{\mathbf{R}}]^{-1}\hat{\mathbf{y}}$. (40)
The reverse process of the diffusion DA thus becomes
$\text{d}\mathbf{v}=-\frac{2}{t}\left[\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\hat{\mathbf{R}}]^{-1}\hat{\mathbf{y}}-\mathbf{v}\right]\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}.$ (41)
Note that the “effective” observation error covariance $\hat{\mathbf{R}}$ implies that, early in the reverse process ($t^{2}\gg r$), the state is drawn towards the observations, $\mathbf{y}_{k}$, but when $t^{2}\ll r$, the state is drawn towards $\mathbf{v}$. Similarly, note that the noise term vanishes as $t\rightarrow 0$ but the drift term takes on a greater significance. These two effects ensure that the ensemble obtained from the diffusion model has the correct mean and variance.
As before, the simple observation operator $\mathbf{H}$ and the simple relation between $\mathbf{x}_{k}$ and $\mathbf{v}$ in (36) imply that all matrices in the reverse process (41) are diagonal. We thus can consider one element of the vector $\mathbf{v}$, which we call $v$ for simplicity. For an observed element of $\mathbf{v}$ we have
$\text{d}v=-\frac{2}{t}\Big{[}\frac{\frac{r}{1+r}}{\frac{r}{1+r}+t^{2}}v+\frac{\frac{t^{2}}{1+r}}{\frac{r}{1+r}+t^{2}}y-v\Big{]}\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}.$ (42)
@@
$\begin{split}v(0)=\frac{\frac{r}{1+r}}{\frac{r}{1+r}+T^{2}}v(T)+\frac{r}{(1+r)^{2}}\frac{T^{2}}{\frac{r}{1+r}(\frac{r}{1+r}+T^{2})}y+\frac{r}{1+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}.\end{split}$ (43)
Note that $v(0)$ is the Gaussian random variable of interest and we can compute its mean as
@@
$\langle v(0)\rangle_{T\rightarrow\infty}=\frac{1}{1+r}y,$ (45)
which is the posterior mean we obtained via the Kalman filter in (31).
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{(\frac{r}{1+r})^{2}}{(\frac{r}{1+r}+T^{2})^{2}}(v(T)-\langle v(T)\rangle)^{2}+\Bigg{\langle}\Bigg{(}\frac{r}{1+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\end{split}$ (46)
[…]
$\begin{split}\Bigg{\langle}\Bigg{(}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}=\Bigg{\langle}\Bigg{(}-\int_{0}^{T}\frac{\sqrt{2t}}{\frac{r}{1+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\\ =2\int_{0}^{T}\frac{t}{(\frac{r}{1+r}+t^{2})^{2}}\text{d}t=\frac{T^{2}}{\frac{r}{1+r}(\frac{r}{1+r}+T^{2})}\end{split}$ (47)
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{r^{2}}{(1+r)^{2}}}{(\frac{r}{1+r}+T^{2})^{2}}T^{2}+\frac{r}{1+r}\frac{T^{2}}{\frac{r}{1+r}+T^{2}}\end{split}$ (48)
@@
which is the same variance as the Bayesian posterior using a climatological prior in (32).
Hence, we have now shown that a diffusion DA system that is trained with a long time-series of system states and observations, samples a Bayesian posterior distribution with a climatological prior.
4.2.3 Diffusion DA (approximate denoiser)
We use the approximate score in equation (22) to develop a diffusion model and then examine the resulting distributions to see how close they are to the Bayesian posterior using a climatological prior. By making use of the same procedure that determined equation (40) we find that the unconditional climatological denoiser is
$D=\frac{1}{1+t^{2}}\mathbf{\mathbf{v}}.$ (50)
Additionally, we have that
[…]
and combining these two equations with (33) leads to the diffusion model
$d\mathbf{v}=-2\frac{t}{1+t^{2}}\mathbf{v}\text{d}t-\frac{2}{r}\frac{t}{1+t^{2}}\mathbf{H}^{T}\Big{[}\mathbf{y}-\frac{1}{1+t^{2}}\mathbf{H}\mathbf{\mathbf{v}}\Big{]}\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}.$ (52)
As in the previous section with the exact denoiser we now examine an observed grid point to obtain the following scalar equation
$dv=2t\frac{1+r(1+t^{2})}{r(1+t^{2})^{2}}v\text{d}t-\frac{2}{r}\frac{t}{1+t^{2}}y\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t}.$ (53)
The solution to this equation is derived in Appendix C and is repeated here
$v(0)=\frac{e^{-\frac{1}{r}\frac{T^{2}}{1+T^{2}}}}{1+T^{2}}v(T)+\Big{[}1-e^{-\frac{T^{2}}{r(1+T^{2})}}\Big{]}y+e^{-\frac{1}{r}}\int_{T}^{0}\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}.$ (54)
The mean of the ensemble produced by this diffusion model is
$\langle v(0)\rangle=\frac{e^{-\frac{1}{r}\frac{T^{2}}{1+T^{2}}}}{1+T^{2}}\langle v(T)\rangle+\Big{[}1-e^{-\frac{T^{2}}{r(1+T^{2})}}\Big{]}y$ (55)
which in the limit as $T\rightarrow\infty$ is
$\langle v(0)\rangle=\Big{[}1-e^{-\frac{1}{r}}\Big{]}y.$ (56)
Upon comparing (56) to the true mean (45) one finds that (56) is biased high, which implies that it draws too closely to the observations.
Similarly, we can compute the variance of the ensemble from the diffusion model as
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{e^{-\frac{2}{r}\frac{T^{2}}{1+T^{2}}}}{(1+T^{2})^{2}}(v(T)-\langle v(T)\rangle)^{2}+\Bigg{\langle}\Bigg{(}e^{-\frac{1}{r}}\int_{T}^{0}\sqrt{2t}\frac{e^{\frac{1}{r}\frac{1}{1+t^{2}}}}{1+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}.\end{split}$ (57)
[…]
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{e^{-\frac{2}{r}\frac{T^{2}}{1+T^{2}}}}{(1+T^{2})^{2}}T^{2}+\frac{r}{2}\Big{[}1-e^{-\frac{2}{r}\frac{T^{2}}{1+T^{2}}}\Big{]}.\end{split}$ (58)
Lastly, evaluating this formula in the limit as $T\rightarrow\infty$ gives
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{r}{2}\Big{[}1-e^{-\frac{2}{r}}\Big{]}.\end{split}$ (59)
A careful comparison with (49) reveals that the approximate denoiser leads to a variance that is too small.
Hence, we have now shown that a diffusion model trained with a long time series of past data and with an approximate denoiser approximates random samples from the posterior with a climatological prior. Due to the approximation in the denoiser, the posterior mean is biased too high and the posterior variance is biased too low.
4.3 Diffusion DA and a cycling prior
We now consider a cycling prior that propagates information from one cycle to the next. Ensemble DA targets the corresponding Bayesian posterior distribution using this cycling prior, i.e. equation (1). We will show that a diffusion DA system that samples a Bayesian posterior distribution with a cycling prior requires re-training at each cycle. The reason is that the prior is equivalent to the training dataset for the diffusion model and therefore the training dataset must be regenerated at each cycle consistent with the latest observations and the “flow-of-the-day”.